# Data-Juicer 2.0: Cloud-Scale Adaptive Data Processing for Foundation Models

### v1

@@
Abstract.
The burgeoning field of foundation models necessitates advanced data processing mechanisms capable of harnessing vast valuable data with varied types utilized by these models. Nevertheless, the current landscape presents unique challenges that traditional data processing frameworks cannot handle effectively, especially with multimodal intricacies. In response, we present Data-Juicer 2.0, a new system offering fruitful data processing capabilities backed by over a hundred operators spanning various modalities like text, image, audio, and video. With seamless compatibility and dedicated optimization to popular dataset hubs like Hugging Face and computing engines like Ray, Data-Juicer 2.0 enhances its predecessor in both usability, efficiency, and programmability. It features an easily accessible user interface layer that supports decoupled Python interactions, RESTful APIs, and conversational commands. Alongside this, it contains a core runtime layer optimized for adaptive execution and management across different dataset scales, processing demands, and computational environments, while shielding unnecessary system details. Extensive empirical evaluations demonstrate Data-Juicer 2.0’s remarkable performance and scalability, highlighting its capability to efficiently process tens of billions of data samples with tens of thousands of CPU cores. The system is publicly available, actively maintained, and broadly adopted in diverse research endeavors, practical applications, and real-world products such as Alibaba Cloud PAI.
@@ Abstract.
1. Introduction
Figure 1. The overview of Data-Juicer 2.0.
Data Processing Challenges for Foundation Models. The advent of foundation models capable of processing textual and multimodal information, exemplified by the GPT-series (OpenAI, 2024), has transformed artificial intelligence by harnessing expansive datasets and substantial computational resources (Wang et al., 2023b; Zhao et al., 2023). The efficacy of these models depends on meticulous data collection, cleaning, annotation, and synthesis (Wang et al., 2023b; Qin et al., 2024b; Bai et al., 2024b), essential for optimizing performance, scalability, and robustness across complex applications.
However, current data processing frameworks, whether for big data (Zaharia et al., 2010) or text-only foundation models (Computer, 2023; Soldaini, Luca and Lo, Kyle and Kinney, Rodney and Naik, Aakanksha and Ravichander, Abhilasha and Bhagia, Akshita and Groeneveld, Dirk and Schwenk, Dustin and Magnusson, Ian and Chandu, Khyathi, 2023; Jennings et al., [n.d.]), are inadequate for handling the complexities of multimodal datasets, due to unique challenges as the field rapidly evolves:
Practitioners lack open-source frameworks specifically designed to manage the diverse data types and semantics integral to foundation models. The processing of these datasets extends beyond standard operations to include specialized tasks such as semantic understanding and cross-modal alignment, which are crucial for optimizing the performance and utility of foundation models (Li et al., 2023b; Tong et al., 2024). For example, the inaugural version, Data-Juicer 1.0  [^1] In this paper, references to Data-Juicer 1.0 pertain to the SIGMOD publication (Chen et al., 2024a) and the open-source codes versioned with the v0.1.2 tag., provided a unified data processing framework for language models with approximately 50 text-only operators (OPs) (Chen et al., 2024a). As user demands and applications evolved rapidly, its complexity increased correspondingly. The need for heterogeneous data processing spanning images, audio, and video underscores the necessity for Data-Juicer 2.0 with enhanced usability, calling for an updated version that can provide advanced coordination mechanisms across various modules, OPs, resources, and models utilized in data processing.
Furthermore, foundation models set specific efficiency demands on data pipelines, emphasizing simple yet computationally intensive operations that traditional big data engines struggle to perform (Team, 2024a; Jennings et al., [n.d.]; Lhoest et al., 2021). In Data-Juicer 1.0, the tightly coupled executors and OP functions hindered stability and efficiency optimization when handling complex, large-scale datasets. In scenarios where a defect is discovered late in handling a massive 10TB dataset at the 9.9TB mark, recovery and cost reduction present formidable challenges.
Moreover, there is a shift towards user interfaces that cater to Python-based ecosystems, preferred by foundation model practitioners (Paszke et al., 2019; Abadi et al., 2016; Rasley et al., 2020), raising additional potential efficiency issues with Python. Addressing new practical needs, Data-Juicer 1.0 showed shortcomings in programmability, such as fragmented development workflows across diverse underlying frameworks like Hugging Face Dataset (Lhoest et al., 2021) and Ray Dataset (Moritz et al., 2018). An ideal solution would enable automatic adaptation to different data modalities, processing requirements, and software-hardware environments, ensuring a smooth, user-friendly development and data processing experience.
Overview of the New Architecture. To address these challenges and respond to emerging needs, we present Data-Juicer 2.0 as illustrated in Fig. 1, emphasizing new functionality and adaptability across hierarchical layers:
Data Processing Capability Layer. Data-Juicer 2.0 provides a broader and more advanced array of capabilities that encompass hundreds of OPs, responding to evolving data processing needs across a wide array of foundation model applications. While Data-Juicer 1.0 focused primarily on data cleansing and lightweight algorithmic implementations, Data-Juicer 2.0 expands the scope beyond the text domain to incorporate image, audio, and video modalities with more functionalities. It offers over 80 new OPs for data analysis, cleaning, synthesis, augmentation, and deduplication, utilizing semantic-aware auxiliary models like GPT and advanced algorithms such as load-balanced distributed union-find. It also includes more compositional OP types to enhance expressive capabilities. Additionally, integrating seamlessly with dataset hubs from Hugging Face (Team, 2024f) and ModelScope (Team, 2024g), it is compatible with massive existing multimodal datasets.
User Interface Layer. Offering a versatile suite of interfaces, Data-Juicer 2.0 caters to various user requirements across multiple operational and abstraction levels. Advanced developers seeking extensive customization will benefit from systematically structured and uniformly designed low-level Python APIs. For simplified service deployment and application integration, lightweight RESTful APIs are available, complemented by a web-based graphical interface that facilitates intuitive and rapid data workflow construction via drag-and-drop features. Moreover, leveraging recent advancements in large language models, natural language processing capabilities are introduced to allow users to issue data processing commands in conversational formats, enhancing accessibility and ease of use.
Core Runtime Layer. At Data-Juicer 2.0’s core is a multi-tiered core runtime optimized for flexible and adaptive processing. It is designed and implemented into four main parts:
(1) A unified Data-Juicer-Dataset (DJ-Dataset) class provides user-friendly programming abstractions and standard capability interfaces across diverse dataset classes, including Hugging Face Dataset, Ray Dataset, and Alibaba Cloud’s MaxCompute DataFrame (Cloud, 2024c), suitable for both standalone and cloud-scale environments;
(2) Decoupled Data-Juicer-Operators (DJ-Operators) classes, which deliver foundational primitives with flexible invocation, automatic optimization, and enhanced capabilities for inspection, integration, and testing. Various design principles and patterns are incorporated (Gamma, 1995), such as Single Responsibility, Strategy, Decorator, Abstract Factory, and Template Method;
(3) Control-panel components orchestrate and manage data processing jobs. These components, like the executor, integrate DJ-Dataset, DJ-Operators, monitors, and other modules seamlessly, forming templates and workflows with proven effectiveness. We extend them beyond dataset-specific development to support data-model co-development (Chen et al., 2024c), utilizing cost-effective model training and evaluation signals via an experimental sandbox;
(4) Adapter features are applied to both DJ-Dataset and DJ-Operators to enable automatic adaptation of data processing tasks to variations in computational requirements (e.g., different OP orchestrations and workloads), software environments (various data engines), and hardware configurations (varying memory, CPU cores, and GPUs). Internally, the Dataset Adapter includes features like engine-agnostic compatibility, streaming I/O, and fault tolerance, while the Operator Adapter facilitates automatic workload-aware OP reordering, OP-wise parallelism, and insight mining.
@@ 1. Introduction
Extensive Evaluation and Usages. We conduct thorough experimental analyses to assess the system’s capabilities under varied workloads, such as Filter-Heavy, Mapper-Heavy, Model-Heavy, Multimodal, and Fuzzy-Deduplication scenarios, using datasets ranging from millions to tens of billions of entries. Our findings provide detailed insights and illuminate performance trade-offs across different use cases and resources, involving Ray and Alibaba MaxCompute with up to 100 nodes and 12,800 cores.

### v2

# Data-Juicer 2.0: Cloud-Scale Adaptive Data Processing for and with Foundation Models

@@
Abstract
The burgeoning field of foundation models necessitates advanced data processing mechanisms capable of harnessing vast and valuable data with various types used by these models. Nevertheless, the current landscape presents unique challenges that traditional data processing frameworks struggle to handle effectively, particularly in handling the complexity of multimodal data. In response, we present Data-Juicer 2.0, a data processing system backed by 100+ data processing operators spanning text, image, video, and audio modalities, supporting more critical tasks including data analysis, synthesis, annotation, and foundation model post-training. With seamless compatibility and dedicated optimization for popular dataset hubs like Hugging Face and computing engines like Ray, it improves upon its predecessor in terms of usability, efficiency, and programmability. It features an easily accessible user interface layer that supports decoupled Python interactions, RESTful APIs, and conversational commands. It contains a new runtime layer optimized for adaptive execution and management across varying dataset scales, processing demands, and computational environments, while hiding unnecessary system details. Extensive empirical evaluations demonstrate Data-Juicer 2.0’s remarkable performance and scalability, highlighting its capability to efficiently process TB-level data with 10k+ CPU cores. The system is publicly available and has been widely adopted in diverse research fields and real-world products such as Alibaba Cloud PAI. We actively maintain it and share insights from practical feedback, with the goal of facilitating research and application of next-generation foundation models.
@@ Abstract
1 Introduction
Data Processing Challenges for Foundation Models. Modern foundation models require sophisticated data pipelines to handle multimodal inputs and evolving paradigms for pre-training and post-training. While existing frameworks address specific aspects of text processing together2023redpajama ; DolmaToolkit or traditional big data workloads apache-spark , they lack essential capabilities for contemporary multimodal workflows. Three critical gaps emerge:
Multimodal Processing Limitations: Current tools provide inadequate support for cross-modal alignment and semantic-aware transformations crucial for vision-language models li2024mgm . The transition from text-only systems like the inaugural version, Data-Juicer 1.0.  [^1] In this paper, references to Data-Juicer 1.0 pertain to the SIGMOD publication chen2024dj and the open-source codes versioned with the v0.1.2 Based on it, there are over 100 pull requests (PRs) for updating to Data-Juicer 2.0, including about 40k and 8k lines of Python code addition and deletion, respectively. to multimodal processing introduces architectural challenges in handling heterogeneous data types and inter-modal relationships.
Efficiency-Scalability Tradeoffs: Traditional big data engines dask and Python-based solutions lhoest-etal-2021-datasets struggle with foundation models’ unique computational patterns - simple per-sample operations applied at petabyte scales. This creates reliability risks in large-scale processing scenarios where late-stage errors can invalidate terabytes of computation.
Ecosystem Fragmentation: Disjoint APIs across popular frameworks (Hugging Face Datasets lhoest-etal-2021-datasets , Ray Data ray ) force practitioners into suboptimal workflow choices. The lack of unified abstractions hampers portable optimization and cross-platform execution.
Architecture Overview. Data-Juicer 2.0 addresses these challenges with a layered architecture optimized for foundation models (Fig. 1), enabling multimodal data processing at scale:
Capability Layer: Extends Data-Juicer 1.0’s 50 text-only operators (OPs) for pre-training to 150+ multimodal OPs supporting text/image/video/audio processing and more post-training tasks. New OPs integrate foundation models (e.g., Tongyi-Qwen qwen3 , SDXL podell2023sdxl ) for semantic-aware filtering and cross-modal synthesis while maintaining compatibility with mainstream dataset hubs hf-dataset-hub ; ms-dataset-hub .
Interface Layer: Provides multi-level APIs balancing flexibility and accessibility. Low-level Python APIs enable custom workflows, while RESTful endpoints and a visual editor support rapid prototyping. Novel agent integration allows natural language specification of data pipelines.
Runtime Layer: Introduces four key enhancements: (1) A unified Data-Juicer-Dataset abstraction spanning local to cloud-scale execution environments; (2) Decoupled OPs with automatic optimization through adaptive batching and resource allocation; (3) Control plane supporting fault-tolerant execution and data-model co-development chen2024djsandbox for insight mining; and (4) Adapter components enabling automatic hardware-software configuration and parallelism across diverse deployments.
@@ 1 Introduction
Extensive Evaluation and Usages. We conduct thorough experimental analyses to assess the system’s data processing capabilities under varied workloads, such as filtering-intensive, model-based, and semantically editable multimodal scenarios, using datasets ranging from millions to tens of billions of samples. We provide actionable insights and performance trade-offs across different use cases and resources, involving Ray and Alibaba MaxCompute maxcompute2024 with up to 100 nodes and 12,800 cores.